# Normalizing flows as an enhanced sampling method for atomistic supercooled liquids

### v1

@@ III.4 Population annealing (PA) and reweighting (RW)
Given a set of $R$ configurations, $\{x_{i}\}$, with $i=1,...,R$ taken from the Boltzmann $NVT$ ensemble at temperature $T_{1}$, it is possible to reweight these configurations to perform an equilibrium average at a different temperature $T_{2}$ [74]. To this end, one assigns a new Boltzmann weight $W_{i}=\exp(-(\beta_{1}-\beta_{2})U(x_{i}))$ to each configuration $i$. The ensemble average of an observable $A(x)$ at $T_{2}$ is given by
@@
Reweighting is used extensively for free energy calculations in molecular simulations [75, 76]. The method, however, only works efficiently for small enough temperature steps so that the weights $W_{i}$ remain meaningful.
In PA, a large set of configurations is used to perform small temperature steps to gradually anneal the temperature to the target low-temperature, while reweighting their Boltzmann weights at each step [22, 77, 23, 78, 79, 80]. In practice, the PA algorithm works as follows. We first create $R$ configurations at an initial, high temperature $\beta_{1}=(k_{B}T_{1})^{-1}$ and evaluate the Boltzmann weight of each configuration $x_{i}$ as $W_{i}=\exp(-(\beta_{1}-\beta_{2})U(x_{i}))$. Here, $T_{2}$ should be slightly smaller than $T_{1}$. We then create on average $\tau_{i}$ copies of each configuration $i$, where $\tau_{i}$ is given by
@@
Recently, different schemes were compared to numerically implement Eq. (4[80]. We apply the “systematic resampling” scheme, which was the most efficient for a constant population size $R$. Following resampling, we finally perform $M$ MD steps on each copy at temperature $T_{2}$ to help thermalize the configurations at the new temperature $T_{2}$. This ends the annealing from $T_{1}$ to $T_{2}$. This annealing step $T_{1}\to T_{2}$ is then repeated several times until the final target temperature $T$ is reached. Each annealing step consists in (i) resampling the population (ii) a small number of $M$ MD steps for each configuration.
@@
Different from MD and PT, there are no separate equilibration and sampling procedures within PA. Nevertheless, it is possible to provide an equivalent time dependence to both $\Delta E_{\text{pot}}(t(R))$ and $c_{V}(t(R))$ by following the evolution of the PA performance as a function of the population size $R$ produced at the initial temperature $T_{1}$, since by definition correct sampling is obtained in the limit $R\to\infty$, just as correct sampling is performed in the large time limit for any of the other algorithms. We can then convert the population size $R$ into a computational timescale using the dictionary $t=10^{4}R$, which corresponds to the effective computational time invested into creating the set of $R$ configurations.
@@ IV Sampling by Normalizing Flows
In Sec. III we established benchmarks for enhanced sampling techniques known from physics. This sets the stage for a thorough analysis of the performance of the ML method of normalizing flows (NF) [33, 81]. Since NF are relatively new methods, we first provide a general introduction before describing our implementation and the main results.
@@ IV.1 Continuous normalizing flows (NF)
The general idea of normalizing flows is to learn an invertible mapping between two probability distributions: a prior distribution $\rho_{P}(x)$, from which we can sample easily (Gaussian random numbers, high temperature liquids), and a target distribution, $\rho_{*}(x)$, in our case the Boltzmann distribution [34]. The mapping is in general only approximate, so one needs to reweight the configurations obtained by the NF. A more accurate mapping leads to lower rejection. Normalizing flows found applications in computer vision [82], sampling via Markov Chain Monte Carlo [83, 37, 84], lattice field theories [85] and condensed matter physics [34, 40].
Boltzmann generators are the first application of normalizing flows for sampling of complex systems in condensed matter [34]. Compared to applications such as image generation, a specific property of Boltzmann generators is that the target distribution is known and corresponds to the Boltzmann distribution
$\rho_{*}(x)=Z_{*}^{-1}\exp(-\beta_{*}U(x)),$ (5)
where $Z_{*}$ is a proportionality constant, $\beta_{*}=(k_{B}T_{*})^{-1}$ is the target inverse temperature and $U(x)$ is the energy. The challenge for Boltzmann generators in statistical physics is to efficiently sample from this distribution using the learned NFs. For a general introduction to Boltzmann generators see Refs. [34, 39].
Here, we use equivariant, continuous normalizing flows [86]. This specific NF has the advantage of being equivariant to translations and rotations and thus mirrors the fundamental symmetries of our system. Continuous NFs transform the prior distribution into the target distribution by learning a time- and space-dependent vector field, $v(x_{v}(t),t)$, $t\in[0,1]$, which can be interpreted as force field,
@@
We define the invertible transformation $F,$ as $x_{v}(t=1)\equiv Fx_{0}$ [^1] Usually this transformation is defined as $T$, see Ref. [86], which we avoid due to the importance of the temperature $T$ in the present study.. Importantly for the calculation of the loss function for training, we can evaluate the transformation of the prior probability distribution [86],
@@
The model used for the force field is a sum of pairwise potentials which depend on the distance between particle pairs [86],
@@
with $d_{ij}(t)=|x_{i}(t)-x_{j}(t)|$. The learnable weights $\{w\}$ of the NF are the parameters of the potential field $\tilde{\Phi}(d,t)$, which is parameterized using Gaussian radial basis functions in both distance $d$ and time $t$, as detailed in Ref. [86]. Our implementation is based on the public code bgflow provided by the authors of this publication. In the following, our goal is to transform an easy-to-sample high-temperature distribution, $\rho_{P}$, at inverse temperature $\beta_{P}$ into a low-temperature target distribution.
@@ IV.2 Loss function and training
Normalizing flows can be trained using the Kuhlback-Leibler divergence as minimizable loss function $L$, which quantifies the similarity between the target distributions and the transformed NF distributions, $L=\alpha D_{\text{KL}}({\rho_{F}}||\rho_{*})+(1-\alpha)D_{\text{KL}}(\rho_{*}||\rho_{\bar{F}}).$ We differentiate in $L$ between two different training contributions. The first term [86],
@@
It has been found in many studies that the training can be improved using a small set of low temperature configurations, $\{x_{*}\}$ sampled from $\rho_{*}(x).$ The second term [86],
@@
quantifies the similarity between the transformed low-temperature configurations, $x=\bar{F}x_{*}$, and the prior distribution. This second contribution is also known as “maximum likelihood” training. We also discretize this equation,
$D_{\text{KL}}(\rho_{*}||\rho_{\bar{F}})=\sum_{i}\left[\beta_{P}U(\bar{F}x_{*})-\int_{0}^{1}\text{d}t\text{ div }v(x_{v}^{i}(t),t)\right],$ (13)
to use it for training the NF. We use $\alpha=0.5$, a specific choice which has little effect on the results if slightly varied. We also tested protocols in which $\alpha$ changes during the training procedure as suggested in Ref. [86], but this did not affect the results qualitatively.
Using a set of high- and low temperature configurations, $\{x_{0}\}$ and $\{x_{*}\}$, enables the optimization of the learnable parameters $\{w\}$, which quantify the strength of the force field $v(x(t),t)$. This step is achieved using the above-defined loss function and stochastic gradient decent. The errors are backpropagated using automatic gradient differentiation of the discretized NF in Eq. (6) implemented by PyTorch, which replaces the typical backpropagation known from artificial neural networks [86].
In our case, however, we face the problem that we do not have access to any low temperature configurations, because generating them is the whole purpose of the NF. We bypass this contradiction by using an iterative procedure. We use reweighting to generate some approximate low-temperature configurations. Using these configurations, we train a NF in the first iteration and apply it to create an improved set of low-temperature samples. In a second iteration we then utilize this improved set for the training of a second NF, which finally produces the low-temperature configurations that are analysed below. We can iteratively improve the performance of the NF with this iterative procedure but using more than two iterations did not lead to significant changes. In fact, we have tested that using low-temperature configurations prepared with SMC for training only marginally improved the performance of the NF, which thus confirms the efficiency of the proposed iterative procedure.
@@ IV.4 Results: Sampling efficiency of NF
We first compare NF to conventional MD results. From Fig. 5(a,b), we conclude that the results for NF are much closer to the SMC groundtruth than what is achieved by MD simulations (see Fig. 2). From Fig. 5(b) we conclude that NF produces an equilibrium ensemble down to $T_{\text{NF}}=0.2$, which is significantly smaller than $T_{\text{MD}}=0.3$. Thus, the NF generative modeling approach is indeed an enhanced sampling method, in the sense that it works better than the physical dynamics in sampling low temperature configurations of the glassy system under study. Given published results regarding generative models for atomistic [39] or glass [47] models, this is an interesting result.
@@ IV.5 Analysis of the effective sample size
One practical consequence of such one-shot annealing is the possibility to define an interpretable effective sample size based on Kish’s formula [88],
@@
In Fig. 6, we show the evolution of $R_{\text{eff}}$ evaluated during the gradual annealing employed for PA in Sec. III.4. We observe a much slower decrease when temperature is reduced, with an effective size that remains quite large, $R_{\rm eff}\sim 10^{4}$ when crossing the temperature $T_{\rm PA}$, indicating that the sample size is a poor indicator of adequate sampling in that case. This presumably results from the resampling of the algorithm whereby the samples that eventually dominate the low temperature behavior are replicated more often than others, which introduces strong correlations in the population. These correlations make the use of Kish formula inefficient. Other independent measures have been suggested for PA to test equilibration, but they require more involved analysis [89].
@@ V.1 Quantitative comparison between techniques
For each technique, we have obtained a temperature below which the assigned sampling task, i.e. measuring the specific heat, starts to fail. This allowed us to rank the various techniques. For the particular glass model studied here, we find that SMC is by far the best technique, with $T_{\rm SMC}=0.12$. Then come the three enhanced algorithms, PA, NF and PT with $T_{\rm PA}=0.19$, $T_{\rm NF}=0.2$, $T_{\rm PT}=0.23$, which all perform much better than conventional MD with $T_{\rm MD}=0.3$. For comparison, we recall that the mode-coupling crossover is near $T=0.3$ and the experimental glass transition temperature near $T=0.15$ [56].
This ranking does not easily translate into an actual computational speedup, or efficiency gain, which may depend on the temperature. For each algorithm and each temperature, we showed that the approach to equilibrium of the energy or the convergence timescale for the specific heat can both be recorded to assign a representative sampling timescale. In practice, we use the former and define a timescale $\tau_{c}$ as $\langle\Delta E_{\text{pot}}(\tau_{c})\rangle_{\rm neq}=0.5$.
@@ V.2 Perspectives
Still, the observed performance of NF should encourage further work towards the development of improved techniques. For instance, it would be interesting to study more complex parametrization of the flow than the one we used. Possible candidates are: equivariant coupling flows [90], which combine the efficiency of coupling flows while maintaining equivariance, or equivariant flow matching [91, 92], which uses alternative loss functions for training [93], or approaches based on diffusion models [94, 95, 96].
@@ Appendix A Details on the NF method
We provide more information on the hyperparameters used in the NF method. Our implementation is based on the bgflow library [34, 86] (https://github.com/noegroup/bgflow), which was extended to include periodic boundary conditions and multiple particle types. Thus any detail provided in Ref. [86] similarly applies to the present manuscript.
Most notably, we discretize the differential equation flow in Eq. (6) using just $N_{t}=1$ (first iteration) or $N_{t}=3$ timesteps (second iteration) in a multi-step fourth order Runge-Kutta scheme [86]. This is a strong reduction of the complexity of the normalizing flows, but we have empirically found that larger values of $N_{t}$ do not improve the results. A consequence of this choice is that the transformation of the configurations amounts to quite small displacements $\Delta x\ll\sigma$ within the particle cages, rather than large-scale rearrangements. We have tried intensively to learn more general models, starting from $T\rightarrow\infty$ (uniformly distributed particles), but none of these models was able to propose acceptable configurations for low temperatures and reach accuracies comparable to the results presented in the main text.

### v2

@@ III.4 Population annealing (PA) and reweighting (RW)
Given a set of $R$ configurations, $\{x_{i}\}$, with $i=1,...,R$ taken from the Boltzmann $NVT$ ensemble at temperature $T_{1}$, it is possible to reweight these configurations to perform an equilibrium average at a different temperature $T_{2}$ Ferrenberg and Swendsen (1988). To this end, one assigns a new Boltzmann weight $W_{i}=\exp(-(\beta_{1}-\beta_{2})U(x_{i}))$ to each configuration $i$. The ensemble average of an observable $A(x)$ at $T_{2}$ is given by
@@
Reweighting is used extensively for free energy calculations in molecular simulations Shen and Hamelberg (2008); Miao et al. (2014). The method, however, only works efficiently for small enough temperature steps so that the weights $W_{i}$ remain meaningful.
In PA, a large set of configurations is used to perform small temperature steps to gradually anneal the temperature to the target low-temperature, while reweighting their Boltzmann weights at each step Hukushima and Iba (2003); Tokdar and Kass (2010); Machta (2010); Wang et al. (2015); Amey and Machta (2018b); Gessert et al. (2023). In practice, the PA algorithm works as follows. We first create $R$ configurations at an initial, high temperature $\beta_{1}=(k_{B}T_{1})^{-1}$ and evaluate the Boltzmann weight of each configuration $x_{i}$ as $W_{i}=\exp(-(\beta_{1}-\beta_{2})U(x_{i}))$. Here, $T_{2}$ should be slightly smaller than $T_{1}$. We then create on average $\tau_{i}$ copies of each configuration $i$, where $\tau_{i}$ is given by
@@
Recently, different schemes were compared to numerically implement Eq. (5Gessert et al. (2023). We apply the “systematic resampling” scheme, which was the most efficient for a constant population size $R$. Following resampling, we finally perform $M$ MD steps on each copy at temperature $T_{2}$ to help thermalize the configurations at the new temperature $T_{2}$. This ends the annealing from $T_{1}$ to $T_{2}$. This annealing step $T_{1}\to T_{2}$ is then repeated several times until the final target temperature $T$ is reached. Each annealing step consists in (i) resampling the population (ii) a small number of $M$ MD steps for each configuration. More extended derivations of the technique and algorithms can be found in Ref. Machta (2010).
@@
Different from MD and PT, there are no separate equilibration and sampling procedures within PA. Nevertheless, it is possible to provide an equivalent time dependence to both $\Delta E_{\text{pot}}(t(R))$ and $c_{V}(t(R))$ by following the evolution of the PA performance as a function of the population size $R$ produced at the initial temperature $T_{1}$, because by definition correct sampling is obtained in the limit $R\to\infty$, just as correct sampling is performed in the large time limit for any of the other algorithms. Convergence in the large time or population limit is obvious, since system size and thus energy barriers are finite. However, our problem is to reach convergence in a tractable computational timescale. We can then convert the population size $R$ into a computational timescale using the dictionary $t=10^{4}R$, which corresponds to the effective computational time invested into creating the set of $R$ configurations.
@@ IV Sampling by Normalizing Flows
In Sec. III we established benchmarks for enhanced sampling techniques known from physics. This sets the stage for a thorough analysis of the performance of the ML method of normalizing flows (NF) Rezende and Mohamed (2015); Papamakarios et al. (2021). Since NF are relatively new methods, we first provide a general introduction before describing our implementation and the main results.
@@ IV.1 Continuous normalizing flows (NF)
The general idea of normalizing flows is to learn an invertible mapping between two probability distributions: a prior distribution $\rho_{P}(x)$, from which we can sample easily (Gaussian random numbers, high temperature liquids), and a target distribution, $\rho_{*}(x)$, in our case the Boltzmann distribution Noé et al. (2019). The mapping is in general only approximate, so one needs to reweight the configurations obtained by the NF. A more accurate mapping leads to lower rejection. Normalizing flows found applications in computer vision Dinh et al. (2016), sampling via Markov Chain Monte Carlo Song et al. (2017); Gabrié et al. (2022); Klein et al. (2024), lattice field theories Albergo et al. (2021) and condensed matter physics Noé et al. (2019); van Leeuwen et al. (2023).
Boltzmann generators are the first application of normalizing flows for sampling of complex systems in condensed matter Noé et al. (2019). Compared to applications such as image generation, a specific property of Boltzmann generators is that the target distribution is known and corresponds to the Boltzmann distribution (see Eq. 1). The challenge for Boltzmann generators in statistical physics is to efficiently sample from this distribution using the learned NFs. For a general introduction to Boltzmann generators see Refs. Noé et al. (2019); Coretti et al. (2022).
Here, we use equivariant, continuous normalizing flows Köhler et al. (2020). This specific NF has the advantage of being equivariant to translations, rotations and permutations and thus mirrors the fundamental symmetries of the underlying physical system. In Ref. Köhler et al. (2020) detailed benchmarking compared to discrete NF layers and gradient flows has been performed on a similar problem showing the superiority of the equivariant continuous NF for particle systems over discrete ones. Continuous NFs transform the prior distribution into the target distribution by learning a time- and space-dependent vector field, $v(x_{v}(t),t)$, $t\in[0,1]$, which can be interpreted as force field,
@@
We define the invertible transformation $F,$ as $x_{v}(t=1)\equiv Fx_{0}$ [^1] Usually this transformation is defined as $T$, see Ref. Köhler et al. (2020), which we avoid due to the importance of the temperature $T$ in the present study.. Importantly for the calculation of the loss function for training, we can evaluate the transformation of the prior probability distribution Köhler et al. (2020),
@@
The model used for the force field is a sum of pairwise potentials which depend on the distance between particle pairs Köhler et al. (2020),
@@
with $d_{ij}(t)=|x_{i}(t)-x_{j}(t)|$. The learnable weights $\{w\}$ of the NF are the parameters of the potential field $\tilde{\Phi}(d,t)$, which is parameterized using Gaussian radial basis functions in both distance $d$ and time $t$. The calculation of the divergence terms is numerically exact and stable, as detailed in Ref. Köhler et al. (2020). Our implementation is based on the public code bgflow provided by the authors of this publication. In the following, our goal is to transform an easy-to-sample high-temperature distribution, $\rho_{P}$, at inverse temperature $\beta_{P}$ into a low-temperature target distribution.
@@ IV.2 Loss function and training
Normalizing flows can be trained using the Kuhlback-Leibler divergence as minimizable loss function $L$, which quantifies the similarity between the target distributions and the transformed NF distributions, $L=\alpha D_{\text{KL}}({\rho_{F}}||\rho_{*})+(1-\alpha)D_{\text{KL}}(\rho_{*}||\rho_{\bar{F}}).$ We differentiate in $L$ between two different training contributions. The first term Köhler et al. (2020),
@@
It has been found in many studies that the training can be improved using a small set of low temperature configurations, $\{x_{*}\}$ sampled from $\rho_{*}(x).$ The second term Köhler et al. (2020),
@@
quantifies the similarity between the transformed low-temperature configurations, $x=\bar{F}x_{*}$, and the prior distribution. This second contribution is also known as “maximum likelihood” training. We also discretize this equation to use it for training the NF.,
$D_{\text{KL}}(\rho_{*}||\rho_{\bar{F}})=\sum_{i}\left[\beta_{P}U(\bar{F}x_{*})-\int_{0}^{1}\text{d}t\text{ div }v(x_{v}^{i}(t),t)\right].$ (13)
We have systematically analyzed the optimal value for $\alpha,$ as discussed in detail in App. A. We find that $\alpha=0.5$ leads to the most stable training procedure and the best final result (see Fig. 9). In particular, this figure also highlights the importance of including the maximum-likelihood training. We also tested protocols in which $\alpha$ changes during the training procedure as suggested in Ref. Köhler et al. (2020), but this did not affect the results qualitatively.
Using a set of high- and low temperature configurations, $\{x_{0}\}$ and $\{x_{*}\}$, enables the optimization of the learnable parameters $\{w\}$, which quantify the strength of the force field $v(x(t),t)$. This step is achieved using the above-defined loss function and stochastic gradient decent. The errors are backpropagated using automatic gradient differentiation of the discretized NF in Eq. (6) implemented by PyTorch, which replaces the typical backpropagation known from artificial neural networks Köhler et al. (2020).
In our case, however, we face the problem that we do not have access to any low temperature configurations, because generating them is the whole purpose of the NF. We bypass this contradiction by using an iterative procedure. We use reweighting to generate some approximate low-temperature configurations. Using these configurations, we train a NF in the first iteration and apply it to create an improved set of low-temperature samples. In a second iteration we then utilize this improved set for the training of a second NF, which finally produces the low-temperature configurations that are analysed below. We can iteratively improve the performance of the NF with this iterative procedure but using more than two iterations did not lead to significant changes. In fact, we have tested that using low-temperature configurations prepared with SMC for training only marginally improved the performance of the NF, which thus confirms the efficiency of the proposed iterative procedure. In fact, our approach is similar in spirit to Ref. Midgley et al. (2023a) which also avoids the usage of target configurations by applying annealed importance sampling.
@@ IV.4 Results: Sampling efficiency of NF
We first compare NF to conventional MD results. From Fig. 6(a,b), we conclude that the results for NF are much closer to the SMC groundtruth than what is achieved by MD simulations (see Fig. 3). From Fig. 6(b) we conclude that NF produces an equilibrium ensemble down to $T_{\text{NF}}=0.2$, which is significantly smaller than $T_{\text{MD}}=0.3$. Thus, the NF generative modeling approach is indeed an enhanced sampling method, in the sense that it works better than the physical dynamics in sampling low temperature configurations of the glassy system under study. Given published results regarding generative models for atomistic Coretti et al. (2022) or glass Ciarella et al. (2023) models, this is an interesting result.
@@ IV.5 Analysis of the effective sample size
One practical consequence of such one-shot annealing is the possibility to define an interpretable effective sample size based on Kish’s formula Kish (1965),
@@
In Fig. 7, we show the evolution of $R_{\text{eff}}$ evaluated during the gradual annealing employed for PA in Sec. III.4. We observe a much slower decrease when temperature is reduced, with an effective size that remains quite large, $R_{\rm eff}\sim 10^{4}$ when crossing the temperature $T_{\rm PA}$, indicating that the sample size is a poor indicator of adequate sampling in that case. This presumably results from the resampling of the algorithm whereby the samples that eventually dominate the low temperature behavior are replicated more often than others, which introduces strong correlations in the population. These correlations make the use of Kish formula inefficient. Other independent measures have been suggested for PA to test equilibration, but they require more involved analysis Amey and Machta (2021).
@@ V.1 Quantitative comparison between techniques
For each technique, we have obtained a temperature below which the assigned sampling task, i.e. measuring the specific heat, starts to fail. This allowed us to rank the various techniques. For the particular glass model studied here, we find that SMC is by far the best technique, with $T_{\rm SMC}=0.12$. Then come the three enhanced algorithms, PA, NF and PT with $T_{\rm PA}=0.19$, $T_{\rm NF}=0.2$, $T_{\rm PT}=0.23$, which all perform much better than conventional MD with $T_{\rm MD}=0.3$. For comparison, we recall that the mode-coupling crossover is near $T=0.3$ and the experimental glass transition temperature near $T=0.15$ Jung et al. (2023b).
This ranking does not easily translate into an actual computational speedup, or efficiency gain, which may depend on the temperature. For each algorithm and each temperature, we showed that the approach to equilibrium of the energy or the convergence timescale for the specific heat can both be recorded to assign a representative sampling timescale. In practice, we use the former and define a timescale $\tau_{c}$ as $\langle\Delta E_{\text{pot}}(\tau_{c})\rangle_{\rm neq}=0.5$. As a rule of thumb, a smaller $\tau_{c}$ implies a smaller computational cost and thus improved performance of the technique.
@@ V.2 Perspectives
Still, the observed performance of NF should encourage further work towards the development of improved techniques. For instance, it would be interesting to study more complex parametrization of the flow than the one we used. Possible candidates are: equivariant coupling flows Midgley et al. (2023b), which combine the efficiency of coupling flows while maintaining equivariance, equivariant flow matching Lipman et al. (2023); Klein et al. (2023), which uses alternative loss functions for training Felardos et al. (2023), annealed flow transport Monte Carlo Arbel et al. (2021), or approaches based on diffusion models Xu et al. (2022); Zheng et al. (2023); Shu et al. (2023). Additionally, it might be possible to combine NF layers with intermittent periods of SMC dynamics to create a stochastic normalizing flow as in Ref. Wu et al. (2020).
@@ Appendix A Details on the NF method
We provide more information on the hyperparameters used in the NF method. Our implementation is based on the bgflow library Noé et al. (2019); Köhler et al. (2020) (https://github.com/noegroup/bgflow), which was extended to include periodic boundary conditions and multiple particle types. Thus any detail provided in Ref. Köhler et al. (2020) similarly applies to the present manuscript.
A.1 Hyperparameter
Most notably, we discretize the differential equation flow in Eq. (6) using just $N_{t}=1$ (first iteration) or $N_{t}=3$ timesteps (second iteration) in a multi-step fourth order Runge-Kutta scheme Köhler et al. (2020). This is a strong reduction of the complexity of the normalizing flows, but we have empirically found that larger values of $N_{t}$ do not improve the results. A consequence of this choice is that the transformation of the configurations amounts to quite small displacements $\Delta x\ll\sigma$ within the particle cages, rather than large-scale rearrangements. We have tried intensively to learn more general models, starting from $T\rightarrow\infty$ (uniformly distributed particles), but none of these models was able to propose acceptable configurations for low temperatures and reach accuracies comparable to the results presented in the main text.